Main INDEX
Monthly INDEX
PREV
NEXT
User name paschke
Log entry time 22:41:52 on May23,2004
Entry number 122547
keyword=paritial run summary, swing shift
-----------------
BeO target
-----------------
Run
1061 BeO target, harp scan 1961
1062 BeO target, dominated by singles
1063 BeO target, dominated by singles
1064 BeO target, dominated by singles
1065 BeO target, dominated by singles
1066 BeO target, dominated by singles
1067 BeO target, harp scan 1962
1068 BeO target, raster on 1.8mmx1.5mm, dominated by singles
1069 BeO target, raster on 2mm x 2mm, dominated by singles
1070 BeO target, DAQ error
17:10 turned on raster, started checking spot size, All BeO target
Run 1068 got 1.8mm x 1.5mm. (Mag currents 10.5 A x 10.6 A)
Run 1069. Requested 2.2x2.5, got 2mm x 2mm (Mag currents 11.6 A x 13.2A)
Run 1070, Changed prescales to emphasize coincidence data. DAQ hang.
Run 1071, new prescales
Run 1072, Final raster setting. about 1.9mm x 2 mm
-----------------
LH2 15 cm target
-----------------
1073 First Data on LH2! 5uA. No visible change in target parameters,
small change in septum temperature.
Changed prescale to match suggested trigger rates at 25 uA.
1074 new prescale settings.
1075 9 uA. RICH HV = 2050V, changed to 2000 V after this run.
1076 RICH HV = 2000 V.
1077 RICH HV = 2000 V.
-----------------
LH2 4 cm target
-----------------
1078 RICH HV=2000 V. Ramping to 30 uA.
1079 RICH HV=2000 V. 30 uA.
1080 RICH HV=2000 V. 35 uA.